# Collision-Free Robot Navigation in Crowded Environments using Learning based Convex Model Predictive Control

### 03 Mar 2024 v1

@@ IV-D Evaluation
Our method demonstrates a higher navigation success rate than the TEB method during all three Stages, outperforming TEB by 10.0%, 12.9%, and 10.7% respectively, which indicates our approach’s superior obstacle avoidance capability in dense dynamic environments. Additionally, our method also reduces navigation time by 1.5s, 1.1s, and 0.7s, and increases speed by 0.46m/s, 0.32m/s, and 0.29m/s, respectively. To analyze the navigation behavior differences between our method and TEB, scenarios from each stage are visualized in Figure 8. While TEB struggles with predicting the movement of dynamic obstacles, leading to potential collisions, our method proactively identifies and navigates through available gaps, allowing for more efficient obstacle avoidance.
@@
Stage Method Success Rate Time (s) Distance (m) Speed (m/s)
5 TEB 77.1% 4.4 4.64 1.11
Ours 87.1% 2.9 4.59 1.57
6 TEB 65.2% 4.0 4.21 1.10
Ours 78.1% 2.9 4.12 1.42
7 TEB 57.8% 3.8 3.87 1.07
Ours 68.5% 3.1 4.22 1.36
@@ IV-E Qualitative Analysis
Figure 9: Navigation Process of Our Method in a Test Scenario at Stage 5 for Qualitative Analysis: Fig. 9 and the first item of Fig. 8 depict the robot’s navigation process during the same test scenario in Stage 5. However, for clearer observation, Fig. 9 omits the radius and timing of dynamic obstacles, as well as the robot’s timing data. Instead, it introduces the convex static obstacle-free regions (indicated by black dashed lines) and a portion of reference trajectories (indicated by green dashed lines). The reference trajectory is formed by connecting long-term and short-term reference points, as detailed in Section III-C.
Figure 9 illustrates the navigation process, reflecting the iterative trajectory optimization procedure depicted in Figure 3. The robot updates the convex static obstacle-free region with the latest LiDAR data. Subsequently, long-term and short-term reference points are sampled within this convex region to form the reference trajectory. The MPC then tracks this reference trajectory to solve the local navigation problem, until the goal endpoint is within the latest convex region and is directly adopted as the long-term reference point. The process ensures static safety by keeping both the reference trajectory and robot motion within the overlapping convex regions. Additionally, the frequent updating of reference points significantly enhances the dynamic safety, ensuring a robust response to environmental changes.
@@ V CONCLUSIONS
This paper introduces a DRL-based navigation framework, which consists of design of action and state spaces, reward function and network architecture. In terms of action space design, our method integrates lidar-generated convex static obstacle-free region to formulate action space that includes both short-term and long-term reference points. For the reward function, we devise a composite reward function enriched with intermediate rewards. We conduct training and experiments across seven Stages. The experimental results conclusively demonstrate that the proposed method significantly enhances robotic performance in both static and dynamic settings, notably excelling in dynamic and crowded environment.

### 14 Mar 2024 v2

@@ IV-D Evaluation
Stage Method Success Rate (%) Time (s) Distance (m) Speed (m/s)
5 TEB[28] 77.1 4.4 4.64 1.11
Ours 87.1 2.9 4.59 1.57
6 TEB 65.2 4.0 4.21 1.10
Ours 78.1 2.9 4.12 1.42
7 TEB 57.8 3.8 3.87 1.07
Ours 68.5 3.1 4.22 1.36
Our method demonstrates higher navigation success rates than the TEB method across all three Stages, outperforming TEB by 10.0%, 12.9%, and 10.7%, respectively. The results highlight the ability of our method to avoid obstacles in highly dynamic environments.
In order to compare the navigation behavior of our method with TEB, several scenarios from each stage are visualized in Figure 6. While TEB struggles with predicting the movement of dynamic obstacles, leading to potential collisions. In contrast, our method moves through available gaps, allowing for more efficient obstacle avoidance.
@@ IV-E Qualitative Analysis
Figure 7: Navigation Process of Our Method in a Test Scenario at Stage 6 for Qualitative Analysis: Omitting the radius and timing of dynamic obstacles, as well as the robot’s timing data. The convex obstacle-free regions are indicated by black dashed lines. And a portion of reference trajectories are indicated by green dashed lines. The reference trajectory is formed by connecting long-term and short-term reference points, as detailed in Section III-C.
Figure 7 illustrates the navigation process, reflecting the iterative trajectory optimization procedure depicted in Figure 3. The robot updates the convex obstacle-free region with the latest LiDAR data. Subsequently, long-term and short-term reference points are sampled within this convex region to form the reference trajectory. The MPC then tracks this reference trajectory to solve the local navigation problem, until the goal endpoint is within the latest convex region and is directly adopted as the long-term reference point. The process aims to promote safety by constraining both the reference trajectory and robot motion within the overlapping convex regions. Additionally, the frequent updating of reference points significantly enhances dynamic safety, ensuring a robust response to environmental changes.
@@ V CONCLUSIONS
This paper introduces a DRL-based navigation framework, which consists of the design of action and state spaces, reward function and network architecture. In terms of action space design, our method integrates lidar-generated convex obstacle-free region to formulate the action space that includes both short-term and long-term reference points. For the reward function, we devise a composite reward function enriched with intermediate rewards. We conduct training and experiments across seven Stages. The experimental results conclusively demonstrate that the proposed method significantly enhances robotic performance in both static and dynamic environments, notably excelling in dynamic and crowded environments.